OpenStudy (anonymous):

PROBLEM SOLVING USING SYSTEMS OF EQUATIONS Donna has half as many purses as Peggy. There are 66 total. How many purses does each girl have?

OpenStudy (anonymous):

x= donna y=peggy x+y= 66 1/2y=x that is all i have so far

OpenStudy (lgbasallote):

let x = no. of purses Peggy has y = no. of purses Donna has y = x/2 x +y = 66 first simplify the first equation... -x + 2y = 0 now perform the elimination thing... 3y = 66 y = 22 Donna has 22 purses Peggy has 44 purses

OpenStudy (anonymous):

thankyou so much, you make it so simple for me!